Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 54     Entry time: 15 Mar 2024, 23:43
Author: Chan Kim 
Group: SPB 
Subject: Shift summary 

**X-ray delivery**

- 6 keV, 3.2 mJ/pulse, stable delivery

**Optical laser delivery**

- 800nm PP laser for jet illumination, stable delivery

**Achievements/Observations**

- NKB focus checked and IUT_Slits optimized
- AGIPD quadrants position finalized
- Measured a lot of water BG runs and few runs with Au nanoparticles (sphere with ~50 nm in diameter)

**Issues**

- SideMic data not saved properly (empty array).
- Found relatively high BG when nozzle is IN since the distance in between nozzle tip & the interaction region is ~100um (even 80 um in some cases).
- DA issue (DAQ was stuck in changing) although all the DA looked healthy. DOC fixed the issue (temporarly) together with Egor, Florian, and Steffen.
- Runs callibration were not working

ELOG V3.1.4-7c3fd00